OpenStudy (anonymous):

[5 -4]+[4 5] perform the indicated operations. if the matrix doesn't exist write impossible

terenzreignz (terenzreignz):

This is just adding components :) You have two 1 by 2 matrices, it results in a new 1 by 2 matrix, where the left number is the sum of the left numbers, and the right number is the sum of the two right numbers. In symbols, \[\huge \left[\begin{matrix}a & b \end{matrix}\right]+\left[\begin{matrix}m & n \end{matrix}\right]=\left[\begin{matrix}a+m & b+n \end{matrix}\right]\]

OpenStudy (anonymous):

so its impossible or possible

terenzreignz (terenzreignz):

Of course it's possible :) You can ALWAYS add two matrices IF they have the same number of rows AND the same number of columns. These two matrices both have only one row, and two columns, so they may be added.

OpenStudy (anonymous):

ty how can we solve that 3[9 4 -3]

terenzreignz (terenzreignz):

This is called "scalar multiplication" It's just like distribution. You simply "bring in" that 3, by creating a new matrix, using the old one, except all the elements inside get multiplied by 3.

OpenStudy (anonymous):

what will be its answer?

terenzreignz (terenzreignz):

Come on... work this out... Remember, multiply 3 to each and every one of the numbers inside the matrix... :)

OpenStudy (anonymous):

[27 12 -9] it woulb the answer

terenzreignz (terenzreignz):

So once again, in general \[\large m \left[\begin{matrix}a & b \\ c & d\end{matrix}\right]=\left[\begin{matrix}ma & mb \\ mc & md\end{matrix}\right]\]

terenzreignz (terenzreignz):

And you are correct :)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

i want to ask one thing more?

terenzreignz (terenzreignz):

sure.

OpenStudy (anonymous):

determine whether the matrices in each pair are inverses x=[1 0 y= [-1 0 1 1 ] 1 1]

terenzreignz (terenzreignz):

You know how to multiply matrices?

OpenStudy (anonymous):

we have to find inverse of matrice

terenzreignz (terenzreignz):

Don't have to. Evaluate this product. \[\large \left[\begin{matrix}1 & 0 \\ 1 & 1\end{matrix}\right]\left[\begin{matrix}-1 & 0 \\ 1 & 1\end{matrix}\right]\]

hartnn (hartnn):

if that product is identity matrix [1 0 0 1 ] then they both are inverses of each other.